OpenStudy (anonymous):

find the solution set x^2-25=0 , separate with a comma ..

OpenStudy (anonymous):

{-5, 5}

OpenStudy (danielbarriosr1):

Yes

OpenStudy (danielbarriosr1):

x^2 -25=0 x^2=25 x= \[\sqrt{25}\] x=5 and -5